How do you solve the following quadratic equations:
-6(u+5)^2=120

3s^2-1=7s^2

Found 2 solutions by stanbon, edjones:
Answer by stanbon(48546)

How do you solve the following quadratic equations:
-6(u+5)^2=120
---
(u+5)^2 = -20
u+5 = 2isqrt(5) or -2isqrt(5)
u = -5+2isqrt(5) or u = -5-2isqrt(5)
=========================================

3s^2-1=7s^2
4s^2 = -1
s^2 = -1/4
s = i/2 or s = -i/2
=======================
Cheers,
Stan H.

Answer by edjones(7311)

-6(u+5)^2=120
(u+5)^2=-20
u^2+10u+25=-20
u^2+10u+45=0
u^2+10u =-45
u^2+10u+25=-45+25 Completing the square
(u+5)^2=-20
u+5=+-sqrt(-20) Take sqrt of each side.
u=-5+-2sqrt(-5)
u=-5+2sqrt(5)i, u=-5-2sqrt(5)i
.
Ed